A certain reproductive element decays according to A(t) = A(0,e^-0.032t). Where t is years. How long will it take to go from 800 grams to 85 grams? (nearest year)
Answer by stanbon(48535)

A certain reproductive element decays according to A(t) = A(0,e^-0.032t). Where t is years. How long will it take to go from 800 grams to 85 grams? (nearest year)
---------------------------------
A(t) = A(0,e^-0.032t).
85 = 800*e^(-0.032t)
e^(-0.032t) = 0.10625
Take the natural log of both sides to get:
-0.032t = -2.24196
t = 70 years (nearest year)